# Efficient Parallel Implementation of the Multiplicative Weight Update Method for Graph-based Linear Programs

### v1

@@ 1. Introduction
Due to this nature, combinatorial algorithms may not scale well on large graphs or large diameter graphs. In addition, some parallelization techniques do not easily generalize between combinatorial algorithms so they don’t naturally allow code reuse. Finally, for certain graph problems, like densest subgraph  (Charikar, 2000), a combinatorial approach does not lend algorithms that are efficient, simple, and almost exact.
@@ 2.1. Positive LP Solvers
Empirical studies of these fast approximate algorithms have been mainly limited to relatively small problems. Koufogiannakis and Young adapt a sequential mixed packing and covering LP and show it outperforms simplex on randomly generated binary matrices with dimension up to $2$ (Koufogiannakis and Young, 2014). Allen-Zhu and Orecchia compare their $\tilde{O}(\epsilon^{-1})$-dependent sequential algorithm (Allen-Zhu and Orecchia, 2014) to the algorithms of Luby and Nisan (Luby and Nisan, 1993) and Awerbuch and Khandekar (Awerbuch and Khandekar, 2009) for solving pure packing LPs with a randomly generated matrix of size $60\times 40$. Jelic et. al. implement a parallel primal-dual method on GPUs to solve positive LPs, although their constraint matrices are randomly generated binary matrices with dimensions up to 25000 (Jelić et al., 2015). The most closely related work to ours is that of Makari et. al. (Makari et al., 2013), who implement a gradient descent algorithm to solve generalized matching on large real-world and synthetic graphs. Their implementation, based on the algorithm from (Awerbuch and Khandekar, 2009), has a $\tilde{O}(1/\epsilon^{5})$ number of iterations, and they confine their attention to a single graph problem. In this paper we focus on only obtaining fractional solutions to the LP problems we consider. The work in  (Makari et al., 2013) also has a rounding step to convert the fractional solution to an integral matching.
@@ 2.2. The MWU Algorithm
For example, to solve a pure packing LP, we embed the objective function as the added constraint, $\frac{1}{M}\mathbb{1}^{T}\bm{x}\geq 1$, where $M$ is the estimate of the maximum value. Then we do binary search over $M$, using Algorithm 1 to determine if the resulting mixed packing and covering LP is a feasible. Since there is one covering constraint, then $\mathrm{smin}_{\eta}(\bm{Cx})=\min(\bm{Cx})$. This exact approximation permits one to scale the step direction (Line 7) by a factor of $2$ in the theoretical analysis, which improves the number of iterations by a factor of $\epsilon$ (Mahoney et al., 2016). Also, noting $\bm{C}=\frac{1}{M}\mathbb{1}^{T}$ and $\nabla\mathrm{smin}_{\eta}(\bm{Cx})=1$, we have $\bm{h}=\frac{1}{M}\mathbb{1}$ (Line 6), so we do not need to explicitly compute $\bm{h}$. Solving a pure covering LP is done via a similar transformation. Solving a mixed covering and packing optimization problem, likewise, involves embedding the constraint that corresponds to the direction of optimization.
@@ 3. Graph Problems as Positive LPs
We now consider several graph problems and LP formulations for them. We first define integer programming (IP) formulations which exactly model the underlying graph problem. We then obtain an LP by relaxing the integrality constraints. Note that for some problems, such as bipartite matching and densest subgraph, the solution to the LP relaxation matches the IP’s solution (i.e., the solution is integral), whereas for NP-hard problems dominating set and vertex cover, there is an integrality gap between the solution to the LP relaxation and IP. We refer the reader (Vazirani, 2013; Williamson and Shmoys, 2011) for the role of LP relaxations in the development of approximation algorithms, and also  (Li et al., 2020) for a performance study on recovering an integral output from a relaxed solution for dominating set as a case study. In this paper, we do not consider rounding for LPs with integrality gaps in both our implementation and in performance comparisons.
@@ 4. MWU with Line Search
We now introduce algorithmic modifications to Algorithm 1 that improve its performance in practice, which we review in 7.3. We write it in Algorithm 2 and mark the modifications in red. Similar to Algorithm 1, we can adapt this algorithm to solve pure packing (e.g., maximum matching) and pure covering (e.g., dominating set, vertex cover) LPs more efficiently, as described in Section 2.2.
@@ 5.1.2. Implicit Representations
In each iteration of MWU, we perform one SpMV with the constraint matrix and with its transpose. As seen in Section 3, the constraint matrices of many graph-based LPs are the vertex-edge incidence matrix of the graph. We notice that the edge $(u,e)$ encoded in an incidence matrix can be described over over a vertex and a vertex pair $(u,(v,w))$ where the pair $(v,w)$ represents an edge in the adjacency matrix. In addition, in the incidence matrix, the value is $1$ if $u=v$ or $u=w$ and 00 otherwise. Therefore, we can store the nonzero data in an incidence matrix implicitly as an adjacency matrix in memory. This reduces the memory cost of the constraint matrix by about half and therefore also reduce the number of accesses to the memory subsystem, particularly cache.
@@ 6.2. Implementations
MWU Implementations. We implement two different versions of the MWU framework presented in Algorithm 2: (1) MWU-PETSc, and (2) MWU-opt. MWU-PETSc relies on an efficient parallel BLAS library PETSc (Balay et al., 2019) while MWU-opt is our hand-optimized implementation using optimizations discussed in Section 5. One modification in our implementation of Algorithm 2 is we do not drop satisfied constraints (i.e., we skip Line 16). This simplifies the implementation, and we did not find this affects convergence to the optimal solution.
PETSc (Balay et al., 2019) is an efficient distributed sparse BLAS library with a Python interface (petsc4py). On Stampede2, we use PETSc with MKL version 19.1.1. We use C++ for our MWU-opt optimized implementation and OpenMP for parallelism. We compile our code with Intel compiler version 19.1.1 and enable -O3, and -mAVX512 compiler flags.
We run the MWU-opt implementation by binding threads to physical cores using numactl –physcpubind. For MWU-PETSc, we find that creating $N$ processes each with 1 thread gives the best performance.
MWU has two parameters: $\epsilon$ and max_iter. $\epsilon$ controls the accuracy of the solution, while max_iter controls the maximum number of MWU iterations, where an LP requiring more than max_iters MWU iterations is deemed infeasible. We set $\epsilon=0.1$ and max_iter=5000. The choice of max_iter is ad-hoc and based on observations that MWU finds $(1+\epsilon)$-relative solutions in at most 1000 iterations across most LP problems and inputs. To verify correctness, we compare the solution from MWU with an exact solution, if available, and report instances where the relative error is greater than $\epsilon$. We implement all applications discussed in Section 3 with MWU.
@@
Specialized Algorithms. We consider optimized custom implementations as baselines for the two implicitly integral LP problems: bipartite matching and densest subgraph. For the former, we use ms-bfs-graft (Azad et al., 2016), which employs the serial Karp-Singer greedy initialization step followed by a specialized breadth-first searches to find augmenting paths. For the latter problem, we used the Graph Based Benchmark Suite’s (Dhulipala et al., 2020) (GBBS) approximate densest subgraph algorithm, which implements Charikar’s greedy 2-approximation algorithm (Charikar, 2000). Both ms-bfs-graft and GBBS are implemented in C++ with OpenMP. We compile ms-bfs-graft using OpenMP and the Intel compiler (version 19.1.1) with the -O2 flag. We compiled GBBS with the g++ compiler version 9.1.0.
@@ 7.1. Comparison of MWU to Other Algorithms
Comparison with Custom Implementations. We compare MWU-opt performance to ms-bfs-graft for bmatch and to GBBS for dense-sub. The MS-BFS algorithm returns an exact solution. MWU-opt returns an approximate solution with $\epsilon=0.1$. The MS-BFS algorithm  (Azad et al., 2016) initializes with a serial Karp-Singer greedy step and finds augmenting paths in parallel using specialized BFS. The performance of ms-bfs-graft heavily depends on the graph structure. In general, we observed the MWU-opt can outperform ms-bfs-graft for graphs with planar structures by 1.8-22.4x for the rgg graphs and usroads. On the other hand, for graphs which contain a strong community-structure or vertices with high degrees, ms-bfs-graft outperforms MWU. For example, amongst the kron graphs, ms-bfs-graft can be up to 450x faster than MWU-opt. For these problems, bmatch general spends much less time on the grafting process than with planar-structured graphs, which we often find to be the dominating cost of bmatch.
@@ 7.2. Parallel Scalability of MWU
For multi-node results, we execute MWU-opt and MWU-PETSc with 64 MPI processes per node and 1 thread per process (this was the most performant configuration for MWU-PETSc on a single machine). We only run experiments where the total number of processes is square, as that is a requirement for our implicit representation. We run all algorithms with Newton’s optimization for step size search and limit Newton’s methods to 5000 iterations. On distributed memory, except for vertex cover on rgg-24, MWU-opt runs faster than MWU-PETSc at scale. For the distributed problem, we also observe almost linear scaling for all graphs except rgg-24 in MWU-opt. A matrix-vector product on the incidence matrix of banded matrices, like rgg-24, corresponds to a 1D problem, so a 1D parallelization (e.g., row-wise distribution of the matrix) is more efficient than a 2D distribution, which is the layout we use.
@@
Figure 4. Scalability of MWU-opt and MWU-PETSc. Axes are in $log_{2}$ scale. All values are normalized to single-threaded execution of MWU-opt. We omit the results for MWU-PETSc, if the execution time is slower than single-threaded execution of MWU-opt.
@@ 7.4.2. Shared-Memory Performance Optimizations
Figure 5c shows the speedup obtained by our optimized implementation relative to the PETSc-based implementation when executing on a single node. In the rest of this section, we report geometric mean speedups when referring to average speedup across graphs. For dom-set, the speedup is obtained from using a favorable format (CSB) and minimizing memory accesses. Our optimizations accelerate matvec operations by 1.8x on average. Although vec and search operations also get speedups, their contribution to overall performance is smaller. On the other hand, for match, bmatch, and vcover problems, we can observe the benefit of specialized vertex-incidence matrix vector multiplications. In these cases, matvec operations are 3.28x, 5.06x, and 3.49x faster on average, respectively. For dense-sub problem, we also see benefits of vertex-edge pair matrix and interweaved-identity matrix specializations. matvec operations are 4.64x faster on average.
@@ Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0022158.

### v3

@@ 2.1. Positive LP Solvers
Empirical studies of these fast approximate algorithms have been mainly limited to relatively small problems. Koufogiannakis and Young adapt a sequential mixed packing and covering LP and show it outperforms simplex on randomly generated binary matrices with dimension up to $2$ (Koufogiannakis and Young, 2014). Allen-Zhu and Orecchia compare their $\tilde{O}(\epsilon^{-1})$-dependent sequential algorithm (Allen-Zhu and Orecchia, 2014) to the algorithms of Luby and Nisan (Luby and Nisan, 1993) and Awerbuch and Khandekar (Awerbuch and Khandekar, 2009) for solving pure packing LPs with a randomly generated matrix of size $60\times 40$. Jelic et. al. implement a parallel primal-dual method on GPUs to solve positive LPs, although their constraint matrices are randomly generated binary matrices with dimensions up to 25000 (Jelić et al., 2015). The most closely related work to ours is that of Makari et. al. (Makari et al., 2013), who implement a gradient descent algorithm to solve generalized matching on large real-world and synthetic graphs. Their implementation, based on the algorithm from (Awerbuch and Khandekar, 2009), has a $\tilde{O}(1/\epsilon^{5})$ number of iterations, and they confine their attention to a single graph problem.
In this paper we focus on only obtaining fractional solutions to the LP problems. Since all LPs are polynomial time solvable, whereas the discrete formulation of some of the graph problems we consider are NP-hard, this allows us to have a more uniform and fair comparison to prior art (e.g., (Makari et al., 2013), as well as state-of-the-art software for solving LPs like Gurobi).
There also exists rounding techniques (which convert a fractional solution to an integral one) or specialized algorithms to solve or approximately solve the discrete problems, but these methods are specific to the problem and have different levels of parallelism, efficiency, and approximation guarantees. For example, an exact parallel rounding technique for a maximum matching problem is implemented in (Makari et al., 2013), but the run time can be three orders of magnitude longer than solving the LP problem. On the other hand, rounding techniques for dominating set are compared in (Li et al., 2020), but these only return approximately optimal solutions and the rounding is not parallelized. There are also specialized libraries for the aforementioned graph problems (Azad et al., 2016; Dhulipala et al., 2020) (more details can be found in Section 6). But again, these implementations are tailored to the problem at hand, hence it would not be fair to compare the run time against our general-purpose solver in terms of obtaining an integral solution.
@@ 2.2. The MWU Algorithm
For example, to solve a pure packing LP, we embed the objective function as the added constraint, $\frac{1}{M}\mathbb{1}^{T}\bm{x}\geq 1$, where $M$ is the estimate of the maximum value, e.g., $M=\sum\limits_{i=1}^{n}\max_{j:p_{ji}>0}1/p_{ji}$. Then we do binary search over $M$, using Algorithm 1 to determine if the resulting mixed packing and covering LP is a feasible. Since there is one covering constraint, then $\mathrm{smin}_{\eta}(\bm{Cx})=\min(\bm{Cx})$. This exact approximation permits one to scale the step direction (Line 7) by a factor of $2$ in the theoretical analysis, which improves the number of iterations by a factor of $\epsilon$ (Mahoney et al., 2016). Also, noting $\bm{C}=\frac{1}{M}\mathbb{1}^{T}$ and $\nabla\mathrm{smin}_{\eta}(\bm{Cx})=1$, we have $\bm{h}=\frac{1}{M}\mathbb{1}$ (Line 6), so we do not need to explicitly compute $\bm{h}$. Solving a pure covering LP is done via a similar transformation. Solving a mixed covering and packing optimization problem, likewise, involves embedding the constraint that corresponds to the direction of optimization.
@@ 3. Graph Problems as Positive LPs
We now consider several graph problems. We first define integer programming (IP) formulations which exactly model the underlying graph problem. We then obtain an LP by relaxing the integrality constraints. For some problems, such as bipartite matching and densest subgraph, the solution to the LP relaxation matches the IP’s solution (i.e., the solution is integral), whereas for NP-hard problems dominating set and vertex cover there is an integrality gap between the solution to the LP relaxation and IP. See (Vazirani, 2013; Williamson and Shmoys, 2011) for the role of LP relaxations in the development of approximation algorithms, and also (Li et al., 2020) for a performance study on rounding an LP relaxation to an integral solution for dominating set. Our goal is to design a general-purpose solver, and the design and performance of rounding schemes are problem dependent (see the end of subsection 2.1 for further details). Therefore, we do not consider rounding in our implementation nor performance comparisons.
@@ 4. MWU with Line Search
We propose two methods for finding a step size in StepSize (Line 11). To motivate these methods, we cast them as 1-D optimization problems, similar to line search methods for (gradient) descent methods (Nocedal and Wright, 2006). However, instead of finding a step size that minimizes the objective function (Nocedal and Wright, 2006), we take the largest step that ensures a local invariance condition is satisfied.
@@ 4.2. Implementing Line Search
A similar line search, which is a coordinate binary search, was proposed in (Quanrud, 2019, Section 2.8), but there are some important differences compared to our binary search. First, rather than taking a step in the full gradient direction $\bm{d}\in\mathbb{R}^{n}$, the coordinate binary search updates sequentially in each of the $n$ indices. Thus, the binary search can have a critical path of length up to $n$ and is therefore not parallel. Second, the coordinate binary search replaces the difference in the smooth min and smooth max from $\Phi$ and $\Psi$ with their respective first-order approximations, which incurs approximation errors in (16) and can lead to more conservative step sizes. By conducting line search in the full gradient $\bm{d}$ and using the exact difference for $\Phi$ and $\Psi$, our proposed binary search improves upon the previous line search in non-trivial ways and can take more aggressive (i.e., larger) step sizes while maintaining feasibility.
@@ 5.1.2. Implicit Representations
In each iteration of MWU, we perform one SpMV with the constraint matrix and with its transpose. As seen in Section 3, the constraint matrices of many graph-based LPs are the vertex-edge incidence matrix of the graph. We notice that the edge $(u,e)$ encoded in an incidence matrix can be described over over a vertex and a vertex pair $(u,(v,w))$ where the pair $(v,w)$ represents an edge in the adjacency matrix. In addition, in the incidence matrix, the value is $1$ if $u=v$ or $u=w$ and 00 otherwise. Therefore, we can store the nonzero data in an vertex-edge incidence matrix $\bm{M}$ implicitly as an adjacency matrix in memory. This reduces the memory cost of the constraint matrix by about half and therefore also reduce the number of accesses to the memory subsystem, particularly cache.
@@ 6.2. Implementations
MWU Implementations.
We implement two different versions of MWU  1: (1) MWU-PETSc, and (2) MWU-opt. MWU-PETSc relies on an efficient parallel BLAS library PETSc (Balay et al., 2019) while MWU-opt is our hand-optimized implementation using optimizations discussed in Section 5. In our implementation of Algorithm 1, we do not drop satisfied constraints (i.e., we skip Line 13). This simplifies the implementation, and we did not find this affects convergence. We set $\epsilon=0.1$ and terminated the algorithm if it exceeds 5000 iterations. To verify correctness, we compare the solution from MWU with an exact solution, if available. max_iter=5000. These parameters control the accuracy and maximum number of iterations of MWU, respectively, which were found by hand-tuning the algorithm. To verify correctness, we compare the solution from MWU with an exact solution, if available.
PETSc is a suite of data structures and routines for large-scale distributed operations, including vector and sparse matrix operations (which calls (sparse) BLAS under the hood), with a Python interface (petsc4py(Balay et al., 2019). On Stampede2, we use PETSc with MKL version 19.1.1. We use C++ for our MWU-opt optimized implementation and OpenMP for parallelism. We compile our code with Intel compiler version 19.1.1 and enable -O3, and -mAVX512 compiler flags. We run the MWU-opt implementation by binding threads to physical cores using numactl –physcpubind. For MWU-PETSc, we find that creating $N$ processes each with 1 thread gives the best performance.
@@
Specialized Algorithms. We consider optimized custom implementations as baselines for the two implicitly integral LP problems: bipartite matching and densest subgraph. For the former, we use ms-bfs-graft (Azad et al., 2016), which employs the serial Karp-Sipser greedy initialization step (Karp and Sipser, 1981) followed by a specialized breadth-first searches to find augmenting paths. For the latter problem, we used the Graph Based Benchmark Suite’s (Dhulipala et al., 2020) (GBBS) approximate densest subgraph algorithm, which implements Charikar’s greedy 2-approximation algorithm (Charikar, 2000). Both ms-bfs-graft and GBBS are implemented in C++ with OpenMP. We compile ms-bfs-graft using OpenMP and the Intel compiler (version 19.1.1) with the -O2 flag. We compiled GBBS with the g++ compiler version 9.1.0.
@@ 7.1. Comparison of MWU to Other Algorithms
Comparison with Custom Implementations. We compare MWU-opt performance to ms-bfs-graft for bmatch and to GBBS for dense-sub. The MS-BFS algorithm returns an exact solution. MWU-opt returns an approximate solution with $\epsilon=0.1$. The MS-BFS algorithm  (Azad et al., 2016) initializes with a serial Karp-Singer greedy step and finds augmenting paths in parallel using specialized BFS. The performance of ms-bfs-graft heavily depends on the graph structure. In general, we observed the MWU-opt can outperform ms-bfs-graft for graphs with planar structures by 1.8-22.4x for the rgg graphs and usroads. On the other hand, for graphs which contain a strong community-structure or vertices with high degrees, ms-bfs-graft outperforms MWU. For example, amongst the kron graphs, ms-bfs-graft can be up to 450x faster than MWU-opt. For these types of graph instances, bmatch generally spends much less time on the grafting process than with planar-structured graphs, which we often find to be the dominating cost of bmatch.
@@ 7.2. Parallel Scalability of MWU
For multi-node results, we execute MWU-opt and MWU-PETSc with 64 MPI processes per node and 1 thread per process (this was the most performant configuration for MWU-PETSc on a single machine). We only run experiments where the total number of processes is square, as that is a requirement for our implicit representation. We run all algorithms with Newton’s optimization for step size search and limit Newton’s methods to 5000 iterations. On distributed memory, except for vertex cover on rgg-24, MWU-opt runs faster than MWU-PETSc at scale. For the distributed problem, we also observe almost linear scaling for all graphs except rgg-24 in MWU-opt. A matrix-vector product on the incidence matrix of banded matrices, like rgg-24, reduces communication on a 1D data distribution pattern, so a 1D parallelization (e.g., row-wise distribution of the matrix) is more efficient than a 2D distribution, which is the layout we use.
@@
Figure 4. Scalability of MWU-opt and MWU-PETSc (see subfigure (j) for the legend). Axes are in $log_{2}$ scale. All values are normalized to single-threaded execution of MWU-opt. We omit the results for MWU-PETSc, if the execution time is slower than single-threaded execution of MWU-opt.
@@ 7.4.2. Shared-Memory Performance Optimizations
Figure 5c shows the speedup obtained by our optimized implementation relative to the PETSc-based implementation when executing on a single node. In this section, we report geometric mean speedups when referring to average speedup across graphs. For dom-set, the speedup is obtained from using a favorable format (CSB) and minimizing memory accesses. Our optimizations accelerate matvec operations by 1.8x on average. Although vec and search operations also get speedups, their contribution to overall performance is smaller. On the other hand, for match, bmatch, and vcover problems, we can observe the benefit of specialized vertex-incidence matrix vector multiplications. In these cases, matvec operations are 3.28x, 5.06x, and 3.49x faster on average, respectively. For dense-sub problem, we also see benefits of vertex-edge pair matrix and interweaved-identity matrix specializations. matvec operations are 4.64x faster on average.
@@ Acknowledgements.
This research has been supported by funding from the United States National Science Foundation (NSF) via grant #1942995 and #2016136, as well as NSF grant CCF-1910149. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0022158.